OpenStudy (anonymous):

Let f(x) = x2 + 3x – 4 and g(x) = x + 5. Find f(x) • g(x). Answer: A) x3 + 5x2 + 14x – 20 B) x3 + 9x2 + 19x – 20 C) x3 + 3x2 + 16x – 20 D) x3 + 8x2 + 11x – 20 ??:)))

OpenStudy (anonymous):

multiply them together, don't forget the distributive property \[(x^2+3x-4)(x+5)\] is our job

OpenStudy (anonymous):

can you do this or do you need help or do you simply need the answer?

OpenStudy (anonymous):

i do need help, but i am just not understanding it. so i don't waste your time anymore trying to get me to understand an answer would at least help me get through the class.:/ im sorryy

OpenStudy (anonymous):

ok first step is \[(x^2+3x-4)(x+5)\] the we multiply. first multiply everything in the first parentheses by x, then by 5 then combine like terms \[(x^2+3x-4)(x+5)=x^3+3x^2-4x+5x^2+15x-20=x^3+8x^2+11x-20\]

OpenStudy (anonymous):

\[x^3+8x^2+11x-20\]

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

yw